HelpMEH123:

Maya has $15 to spend. She spent $5.50, including tax, to buy a notebook. She needs to save $7.75, but she wants to buy a snack. If cookies cost $0.25 per package including tax, what inequality would show the maximum number of packages that Maya can buy? Solve your inequality by showing your work and steps.

HelpMEH123:

HELP MEH

XioGonz:

1 attachment
jimthompson5910:

"Maya has $15 to spend" and "She needs to save $7.75" This means she can spend at most 15-7.75 = 7.25 dollars. If she spends any higher, then it will eat into the amount she wants to save. Now onto the next part " She spent $5.50" and "cookies cost $0.25 per package" This means she spends 5.50 + 0.25x x = number of packages she buys Because she spends 5.50 + 0.25x dollars and the most she can spend is 7.25 dollars, this means we can form the inequality \(\large 5.50 + 0.25x \le 7.25\) I'll let you solve for x

XxAngelGirlxX:

1 attachment
jimthompson5910:

If x = 8, then, \(\large 5.50 + 0.25x \le 7.25\) \(\large 5.50 + 0.25*8 \le 7.25\) \(\large 5.50 + 2 \le 7.25\) \(\large 7.50 \le 7.25\) But the last inequality is false because 7.50 is actually larger than 7.25

jimthompson5910:

Though x = 7 does work because \(\large 5.50 + 0.25x \le 7.25\) \(\large 5.50 + 0.25*7 \le 7.25\) \(\large 5.50 + 1.75 \le 7.25\) \(\large 7.25 \le 7.25\) which is a true inequality

BluSaint:

wat de hek